# Language Inclusion for Boundedly-Ambiguous Vector Addition Systems is Decidable

### v4

@@ 2. Preliminaries
    it is in the “complement“ i.e. for every $i\in\{1,\ldots,n\}$ there is a coordinate $f(i)\in\{1,\ldots,d\}$ such that $x[f(i)]<u_{i}[f(i)]$.
@@
    it is maximal, i.e. for every $y$ that satisfies the first property holds $x\not\preceq y$.
@@
An example language which can be recognised by $k$-ambiguous VASS, but probably cannot be recognised by a $(k-1)$-ambiguous VASS is the following: language of words over $\{a,b\}$ with exactly $k$ letters $b$, which divides the word into $k+1$ blocks consisting of letters $a$, such that the first block is not the shortest one.block it decrements the counter. A $k$-ambiguous $1$-VASS increments the counter on each letter $a$ in the first block, later guesses a block which is the shortest one, and in that guessed block it decrements the counter. As there are at most $k$ other blocks, which may be guessed, there are at most $k$ accepting runs for each word. However, it is not clear how to accept this language by a $(k-1)$-ambiguous VASS. The problematic scenario is if the first block is the longest one.
@@
If $L(V_{1})=L(V_{2})$?
In the sequel, we are mostly interested in VASS with the set of final configurations $F$ of some special form. We extend the order $\preceq$ on the vectors of $\mathbb{N}^{d}$ to configurations from $Q\times\mathbb{N}^{d}$ in a natural way: we say that $q_{1}(v_{1})\preceq q_{2}(v_{2})$ if $q_{1}=q_{2}$ and $v_{1}\preceq v_{2}$. We define the notions of upward-closed, downward-closed, up-atom and down-atom the same as for vectors. As Proposition 6 holds for any well quasi-order, it also applies to $Q\times\mathbb{N}^{d}$. Proposition 7 applies here as well, as the upper bound on the size can be shown separately for each state. Let the set of final configurations of VASS $V$ be $F$. If $F$ is upward-closed then we call $V$ an upward-VASS. If $F$ is closed downward, then we call $V$ a downward-VASS. For two sets $A\subseteq\mathbb{N}^{a}$, $B\subseteq\mathbb{N}^{b}$ and a subset of coordinates $J\subseteq[1,a+b]$ by $A\times_{J}B$ we denote the set of vectors in $\mathbb{N}^{a+b}$ which projected into coordinates in $J$ belong to $A$ and projected into coordinates outside $J$ belong to $B$. If $F=\bigcup_{i\in[1,n]}q_{i}(U_{i}\times_{J_{i}}D_{i})$ where for all $i\in[1,n]$ we have $J_{i}\subseteq[1,d]$, $U_{i}\subseteq\mathbb{N}^{|J_{i}|}$ are up-atoms and $D_{i}\subseteq\mathbb{N}^{d-|J_{i}|}$ are down-atoms then we call $V$ an updown-VASS. In the sequel we write simply $\times$ instead of ×$\times_{J}$, as the set of coordinates $J$ is never relevant. If $F=\{c_{F}\}$ is a singleton, then we call $V$ a singleton-VASS. As in this paper we mostly work with upward-VASS we often say simply a VASS instead of an upward-VASS. In other words, if not indicated otherwise, we assume that the set of final configurations $F$ is upward-closed.
@@ A VASS $V=(\Sigma,Q,T,c_{I},F)$
Does $c_{I}\longrightarrow c_{F}$ in $V$ for some $c_{F}\in F$?
@@ 3.1. Lower bound
For each $d$-dimensional singleton-VASS $V$ with final configuration being $c_{F}=q_{F}(0^{d})$ one can construct in polynomial time two deterministic $(d+1)$-dimensional upward-VASS $V_{1}$ and $V_{2}$ such that
@@ 3.2. Upper bounds
Let $V=(\Sigma,Q,T,c_{I},F)$ be a deterministic $d$-VASS. We aim at constructing a $d$-dimensional downward-VASS $V^{\prime}$ such that $L(V^{\prime})=\overline{L(V)}$. Before constructing $V^{\prime}$ let us observe that there are three possible scenarios for a word $w$ to be not in $L(V)$. The first scenario (1) is that the only run over $w$ in $V$ finishes in a non-accepting configuration. Another possibility is that there is even no run over $w$. Namely, for some prefix $va$ of $w$ where $v\in\Sigma^{*}$ and $a\in\Sigma$ we have $c_{I}\stackrel{{\scriptstyle v}}{{\longrightarrow}}c$ for some configuration $c$ but there is no transition from $c$ over the letter $a$ as either (2) a possible transition over $a$ would decrease some of the counters below zero, (3) there is no such transition possible in $V$ in the state of $c$. For each case we separately design a part of a downward-VASS accepting it. Cases (1) and (3) are simple. For case (2) we nondeterministically guess the moment when the run would go below zero and freeze the configuration at that moment. Then, at the end of the word, we check if our guess was correct. Notice that the set of configurations from which a step labelled with a letter $a$ would take a counter below zero is downward-closed, so we can check the correctness of our guess using a downward-closed accepting condition. A detailed proof follows.
@@
Now we describe the second part of $V^{\prime}$ which is responsible for words rejected by $V$ because of the scenario (2). The idea is to guess when the run over $w$ is finished. For each label $a\in\Sigma$ we add $(p^{\prime},a,0^{d},p^{\prime\prime})$ to $T^{\prime}$ where $p^{\prime}=(p,-)$ and $q^{\prime\prime}=(p,(a,2))$. The idea is that the run reaches the configuration in which the transition labeled with $a$ cannot be fired. Now we have to check that our guess is correct. In the state $(p,(a,2))$ for $t\in T$ no transition changes the configuration. Namely for each $p^{\prime\prime}=(p,(a,2))\in Q\times\Sigma$ and each $b\in\Sigma$ we add to $T^{\prime}$ a transition $(p^{\prime\prime},b,0^{d},p^{\prime\prime})$. We now add to $F^{\prime}$ the set $F_{2}=\{(p,(a,2))(v)\mid v+\textup{eff}(t)\not\in\mathbb{N}^{d}\text{ for all }t=(p,a,\textup{eff}(t),q)\in T\text{ where }q\in Q\}$. Notice that $F_{2}$ can be easily represented as a polynomial union of down-atoms. It is easy to see that indeed $V^{\prime}$ accepts by $F_{2}$ exactly words $w$ such that there is a run of $V$ over some prefix $v$ of $w$ but reading the next letter would decrease one of the counters below zero.
The last part of $V^{\prime}$ is responsible for the words $w$ rejected by $V$ because of the scenario (3), that is, $w$ has a prefix $va$ such that there is a run over $v\in\Sigma^{*}$ in $V$ but then in the state of the reached configuration there is no transition over the letter $a\in\Sigma$. To accept such words for each state $p\in Q$ and letter $a\in\Sigma$ such that there is no transition of a form $(p,a,v,q)\in T$ for any $v\in\mathbb{Z}^{d}$ and $q\in Q$ we add to $T^{\prime}$ transition $((p,-),a,0^{d},(p,(a,3)))$. In each state $p^{\prime}=(p,(a,3))\in Q\times\Sigma$, we have a transition $(p^{\prime},b,0^{d},p^{\prime})$ for each $b\in\Sigma$. We also add to $F^{\prime}$ the set $F_{3}=\{(p,(a,3))(v)\mid v\in\mathbb{N}^{d}\text{ and there is no }(p,a,u,q)\in T\text{ for }u\in\mathbb{Z}^{d}\text{ and }q\in Q\}$. The size of $F_{3}$ is polynomial w.r.t. $T$.
@@
There is an algorithm, that for a deterministic HVASS $V$ computes, in exponential time, a downward-$\varepsilon$-VASS which recognises the complement of the language of $V$.
@@
There is an algorithm that, for a $k$-control-deterministic $d$-VASS $V$ constructs, in exponential time, a $(k\cdot d)$-dimensional downward-VASS which recognises the complement of the language of $V$.
@@
We first sketch our solution. In the construction $(k\cdot d)$-dimensional downward-VASS $V^{\prime}$ simulates $k$ copies of $V$ which take care of at most $k$ different maximal runs of $V$. The accepting condition $F^{\prime}$ of $V^{\prime}$ verifies whether in all copies there is a reason that the simulated maximal runs do not accept. The reasons why each individual copy does not accept are the same as in Theorem 15.
@@ 4. Unambiguous VASS
[Theorem 24 in [CLM+18]] For any two VASS languages $L_{1},L_{2}\subseteq\Sigma^{*}$ if $L_{1}\cap L_{2}=\emptyset$ then $L_{1}$ and $L_{2}$ are regular-separable and one can compute the regular separator in elementary time.
@@
Let $V$ be an unambiguous VASS over $\Sigma$ and $\mathcal{F}$ be a finite family of regular languages separating languages of $V$. Suppose $M$ is a monoid with homomorphism $h:\Sigma^{*}\to M$ recognising every language in $\mathcal{F}$. Then HVASS $V_{(M,h)}$ is deterministic.
@@ 5. Boundedly-ambiguous VASS
Now we consider the control automaton $\mathcal{A}$ of VASS $V^{1}$. We compute a monoid $M$ together with a homomorphism $h:\Sigma^{*}\to M$ that recognises all the state languages of $\mathcal{A}$. Then we construct the automaton $\mathcal{A}_{(M,h)}$. Note that the decoration of a VASS produces an HVASS, but as we decorate an automaton i.e. 00-VASS we get a 00-HVASS which is also a finite automaton. Based on $\mathcal{A}_{(M,h)}$ we construct a VASS $V^{2}$. We add a vector to every transition in $\mathcal{A}_{(M,h)}$ to produce a VASS that recognises the $(M,h)$-decoration of the language of VASS $V^{1}$. Precisely, if we have a transition $((p,m),(a,m^{\prime}),(q,m^{\prime}))$ in $\mathcal{A}_{(M,h)}$ then it is created from the transition $(p,a,q)$ in $\mathcal{A}$, which originates from the transition $(p,a,v,q)$ in $V^{1}$. So in $V^{2}$ we label $((p,m),(a,m^{\prime}),(q,m^{\prime}))$ with $v$ i.e. we have the transition $((p,m),(a,m^{\prime}),v,(q,m^{\prime}))$. Similarly, based on $V^{1}$, we define initial and final configurations in $V^{2}$. It is easy to see that there is a bijection between accepting runs in $V^{1}$ and accepting runs in $V^{2}$. By Lemma 30 $\mathcal{A}_{(M,h)}$ is $k$-deterministic which immediately implies that $V^{2}$ is $k$-control-deterministic as well.
@@ 5.1. Proof of Lemma 29
By Lemma 31 we can compute a threshold $\overline{M}$ such that the $\overline{M}$-abstraction of $\overline{V}$ has no accepting run. We claim now that the $\overline{M}$-abstraction of $V$, namely $V_{\overline{M}}$ is $k$-ambiguous. Assume otherwise; let $V_{\overline{M}}$ have at least $k+1$ runs over some word $w$. Then it is easy to see that the $\overline{M}$-abstraction of $\overline{V}$ has an accepting run over $w$, which is a contradiction. Thus indeed $V_{\overline{M}}$ is $k$-ambiguous and extending the control automaton of $V$ by $M$-bounded counters in a way as in $V_{\overline{M}}$ (such that the control automaton becomes exactly $V_{\overline{M}}$) finishes the proof of Lemma 29
@@ 5.2. Proof of Lemma 30
To prove the second claim, we assume that $w$ is accepted by the automaton $\mathcal{A}$ as otherwise there is just one maximal run of length zero. Let ${X^{{w_{(M,h)}}}_{(M,h)}}=\{(q_{1},m_{1}),(q_{2},m_{1}),\ldots,(q_{\ell},m_{1})\}_{\textsc{mul}}$ be the multiset of states reached by the runs labelled ${w_{(M,h)}}$. Furthermore, because of the proof that all the maximal runs have the same length, we conclude that there are exactly $\ell$ runs along ${w_{(M,h)}}$. Moreover, all of them are accepting, as they end in states with the monoid element $h(\epsilon)$. Observe that the states along the runs of ${w_{(M,h)}}$ never differ on the second component, as the second component is determined by the image $h$ on the suffix of the word. Thus, if two runs over ${w_{(M,h)}}$ differ, then they differ at the first component and in consequence they induce two different runs along $w$ in $\mathcal{A}$. For example, such two different runs over ${w_{(M,h)}}$ with the first difference after the letter $(a_{i+1},m_{i+1})$:

### v5

@@ 2. Preliminaries
    it is in the complement of $U$ i.e. for every $i\in\{1,\ldots,n\}$ there is a coordinate $f(i)\in\{1,\ldots,d\}$ such that $x[f(i)]<u_{i}[f(i)]$.
@@
    it is maximal, i.e. for every $y\neq x$ that satisfies the first property it holds $x\not\preceq y$.
@@
An example language which can be recognised by $k$-ambiguous VASS, but probably cannot be recognised by a $(k-1)$-ambiguous VASS is the following: language of words over $\{a,b\}$ with exactly $k$ letters $b$, which divides the word into $k+1$ blocks consisting of letters $a$, such that the first block is not the shortest one. A $k$-ambiguous $1$-VASS increments the counter on each letter $a$ in the first block, later guesses a block which is the shortest one, and in that guessed block it decrements the counter. As there are at most $k$ other blocks, which may be guessed, there are at most $k$ accepting runs for each word. However, it is not clear how to accept this language by a $(k-1)$-ambiguous VASS. The problematic scenario is if the first block is the longest one.
@@
Does $L(V_{1})=L(V_{2})$ hold?
In the sequel, we are mostly interested in VASS with the set of final configurations $F$ of some special form. We extend the order $\preceq$ on the vectors of $\mathbb{N}^{d}$ to configurations from $Q\times\mathbb{N}^{d}$ in a natural way: we say that $q_{1}(v_{1})\preceq q_{2}(v_{2})$ if $q_{1}=q_{2}$ and $v_{1}\preceq v_{2}$. We define the notions of upward-closed, downward-closed, up-atom and down-atom the same as for vectors. As Proposition 6 holds for any well quasi-order, it also applies to $Q\times\mathbb{N}^{d}$. Proposition 7 applies here as well, as the upper bound on the size can be shown separately for each state. Let the set of final configurations of VASS $V$ be $F$. If $F$ is upward-closed then we call $V$ an upward-VASS. If $F$ is closed downward, then we call $V$ a downward-VASS. For two sets $A\subseteq\mathbb{N}^{a}$, $B\subseteq\mathbb{N}^{b}$ and a subset of coordinates $J\subseteq[1,a+b]$ by $A\times_{J}B$ we denote the set of vectors in $\mathbb{N}^{a+b}$ which projected into coordinates in $J$ belong to $A$ and projected into coordinates outside $J$ belong to $B$. If $F=\bigcup_{i\in[1,n]}q_{i}(U_{i}\times_{J_{i}}D_{i})$ where for all $i\in[1,n]$ we have $J_{i}\subseteq[1,d]$, $U_{i}\subseteq\mathbb{N}^{|J_{i}|}$ are up-atoms and $D_{i}\subseteq\mathbb{N}^{d-|J_{i}|}$ are down-atoms then we call $V$ an updown-VASS. In the sequel we write simply $\times$ instead of ×$\times_{J}$, as the set of coordinates $J$ is never relevant. If $F=\{c_{F}\}$ is a singleton, then we call $V$ a singleton-VASS.
Notation. As in this paper we mostly work with upward-VASS we often say simply a VASS instead of an upward-VASS. In other words, if not indicated otherwise, we assume that the set of final configurations $F$ is upward-closed.
@@ A VASS $V=(\Sigma,Q,T,c_{I},F)$
Is $c_{F}$ reachable from $c_{I}$ in $V$ for some $c_{F}\in F$?
@@ 3.1. Lower bound
For each $d$-dimensional singleton-VASS $V$ with final configuration being $c_{F}=q_{F}(0^{d})$ one can construct in polynomial time two deterministic $(d+1)$-dimensional upward-VASSes $V_{1}$ and $V_{2}$ such that
@@ 3.2. Upper bounds
Let $V=(\Sigma,Q,T,c_{I},F)$ be a deterministic $d$-VASS. We aim at constructing a $d$-dimensional downward-VASS $V^{\prime}$ such that $L(V^{\prime})=\overline{L(V)}$. Before constructing $V^{\prime}$ let us observe that there are three possible scenarios for a word $w$ to be not in $L(V)$. The first scenario (1) is that the only run over $w$ in $V$ finishes in a non-accepting configuration. Another possibility is that there is even no run over $w$. Namely, for some prefix $va$ of $w$ where $v\in\Sigma^{*}$ and $a\in\Sigma$ we have $c_{I}\stackrel{{\scriptstyle v}}{{\longrightarrow}}c$ for some configuration $c$ but there is no transition from $c$ over the letter $a$ as either (2) all possible transitions over $a$ would decrease some of the counters below zero, (3) there is no such transition possible in $V$ in the state of $c$. For each case we separately design a part of a downward-VASS accepting it. Cases (1) and (3) are simple. For case (2) we nondeterministically guess the moment when the run would go below zero and freeze the configuration at that moment. Then, at the end of the word, we check if our guess was correct. Notice that the set of configurations from which a step labelled with a letter $a$ would take a counter below zero is downward-closed, so we can check the correctness of our guess using a downward-closed accepting condition. A detailed proof follows.
@@
Now we describe the second part of $V^{\prime}$ which is responsible for words rejected by $V$ because of the scenario (2). The idea is to guess when the run over $w$ is finished. For each label $a\in\Sigma$ we add $(p^{\prime},a,0^{d},p^{\prime\prime})$ to $T^{\prime}$ where $p^{\prime}=(p,-)$ and $p^{\prime\prime}=(p,(a,2))$. The idea is that the run reaches the configuration in which the transition labeled with $a$ cannot be fired. Now we have to check that our guess is correct. In the state $(p,(a,2))$ for $t\in T$ no transition changes the configuration. Namely for each $p^{\prime\prime}=(p,(a,2))\in Q\times\Sigma$ and each $b\in\Sigma$ we add to $T^{\prime}$ a transition $(p^{\prime\prime},b,0^{d},p^{\prime\prime})$. We now add to $F^{\prime}$ the set $F_{2}=\{(p,(a,2))(v)\mid v+\textup{eff}(t)\not\in\mathbb{N}^{d}\text{ for all }t=(p,a,\textup{eff}(t),q)\in T\text{ where }q\in Q\}$. Notice that $F_{2}$ can be easily represented as a polynomial union of down-atoms. It is easy to see that indeed $V^{\prime}$ accepts by $F_{2}$ exactly words $w$ such that there is a run of $V$ over some prefix $v$ of $w$ but reading the next letter would decrease one of the counters below zero. The last part of $V^{\prime}$ is responsible for the words $w$ rejected by $V$ because of the scenario (3), that is, $w$ has a prefix $va$ such that there is a run over $v\in\Sigma^{*}$ in $V$ but then in the state of the reached configuration there is no transition over the letter $a\in\Sigma$. To accept such words for each state $p\in Q$ and letter $a\in\Sigma$ such that there is no transition of a form $(p,a,v,q)\in T$ for any $v\in\mathbb{Z}^{d}$ and $q\in Q$ we add to $T^{\prime}$ transition $((p,-),a,0^{d},(p,(a,3)))$. In each state $p^{\prime}=(p,(a,3))\in Q\times\Sigma$, we have a transition $(p^{\prime},b,0^{d},p^{\prime})$ for each $b\in\Sigma$. We also add to $F^{\prime}$ the set $F_{3}=\{(p,(a,3))(v)\mid v\in\mathbb{N}^{d}\text{ and there is no }(p,a,u,q)\in T\text{ for }u\in\mathbb{Z}^{d}\text{ and }q\in Q\}$. The size of $F_{3}$ is polynomial w.r.t. $T$.
@@
There is an exponential time algorithm, that for a deterministic HVASS $V$ computes a downward-$\varepsilon$-VASS, which recognises the complement of the language of $V$.
@@
There is an exponential time algorithm, that for a $k$-control-deterministic $d$-VASS $V$ constructs a $(k\cdot d)$-dimensional downward-VASS, which recognises the complement of the language of $V$.
@@
We first sketch our solution. In the algorithm the $(k\cdot d)$-dimensional downward-VASS $V^{\prime}$ simulates $k$ copies of $V$ which take care of at most $k$ different maximal runs of $V$. The accepting condition $F^{\prime}$ of $V^{\prime}$ verifies whether in all copies there is a reason that the simulated maximal runs do not accept. The reasons why each individual copy does not accept are the same as in Theorem 15.
@@ 4. Unambiguous VASS
[Theorem 24 in [CLM+18]] For any two VASS languages $L_{1},L_{2}\subseteq\Sigma^{*}$ if $L_{1}\cap L_{2}=\emptyset$ then $L_{1}$ and $L_{2}$ are regular-separable and one can compute a regular separator in elementary time.
@@
Let $V$ be an unambiguous VASS over $\Sigma$ and $\mathcal{F}$ be a finite family of regular languages separating languages of $V$. Suppose $M$ is a monoid with homomorphism $h:\Sigma^{*}\to M$ recognising every language in $\mathcal{F}$. Then the HVASS $V_{(M,h)}$ is deterministic.
@@ 5. Boundedly-ambiguous VASS
Now we consider the control automaton $\mathcal{A}$ of VASS $V^{1}$. We compute a monoid $M$ together with a homomorphism $h:\Sigma^{*}\to M$ that recognises all the state languages of $\mathcal{A}$. Then we construct the automaton $\mathcal{A}_{(M,h)}$. Note that the decoration of a VASS produces an HVASS, but as we decorate an automaton i.e. 00-VASS we get a 00-HVASS which is also a finite automaton. Based on $\mathcal{A}_{(M,h)}$ we construct a VASS $V^{2}$. We add a vector to every transition in $\mathcal{A}_{(M,h)}$, to produce a VASS, that recognises the $(M,h)$-decoration of the language of VASS $V^{1}$. Precisely, if we have a transition $((p,m),(a,m^{\prime}),(q,m^{\prime}))$ in $\mathcal{A}_{(M,h)}$ then it is created from the transition $(p,a,q)$ in $\mathcal{A}$, which originates from the transition $(p,a,v,q)$ in $V^{1}$. So in $V^{2}$ we label $((p,m),(a,m^{\prime}),(q,m^{\prime}))$ with $v$ i.e. we have the transition $((p,m),(a,m^{\prime}),v,(q,m^{\prime}))$. Similarly, based on $V^{1}$, we define initial and final configurations in $V^{2}$. It is easy to see that there is a bijection between accepting runs in $V^{1}$ and accepting runs in $V^{2}$. By Lemma 30 $\mathcal{A}_{(M,h)}$ is $k$-deterministic which immediately implies that $V^{2}$ is $k$-control-deterministic as well.
@@ 5.1. Proof of Lemma 29
By Lemma 31 we can compute a threshold $\overline{M}$ such that the $\overline{M}$-abstraction of $\overline{V}$ has no accepting run. We claim now that the $\overline{M}$-abstraction of $V$, namely $V_{\overline{M}}$ is $k$-ambiguous. Assume otherwise; let $V_{\overline{M}}$ have at least $k+1$ runs over some word $w$. Then it is easy to see that the $\overline{M}$-abstraction of $\overline{V}$ has an accepting run over $w$, which is a contradiction. Thus indeed $V_{\overline{M}}$ is $k$-ambiguous and extending the control automaton of $V$ by $M$-bounded counters in a way as in $V_{\overline{M}}$ (such that the control automaton becomes exactly $V_{\overline{M}}$) finishes the proof of Lemma 29.
@@ 5.2. Proof of Lemma 30
To prove the second claim, we assume that $w$ is accepted by the automaton $\mathcal{A}$ as otherwise there is just one maximal run of length zero. Let
${X^{{w_{(M,h)}}}_{(M,h)}}=\{(q_{1},m_{1}),(q_{2},m_{1}),\ldots,(q_{\ell},m_{1})\}_{\textsc{mul}}$
be the multiset of states reached by the runs labelled ${w_{(M,h)}}$. Furthermore, because of the proof that all the maximal runs have the same length, we conclude that there are exactly $\ell$ runs along ${w_{(M,h)}}$. Moreover, all of them are accepting, as they end in states with the monoid element $h(\epsilon)$. Observe that the states along the runs of ${w_{(M,h)}}$ never differ on the second component, as the second component is determined by the image $h$ on the suffix of the word. Thus, if two runs over ${w_{(M,h)}}$ differ, then they differ at the first component and in consequence they induce two different runs along $w$ in $\mathcal{A}$. For example, such two different runs over ${w_{(M,h)}}$ with the first difference after the letter $(a_{i+1},m_{i+1})$: